OpenStudy (anonymous):

One of two supplementary angles is seventeen times as large as the other. Find the measure of each angle.

OpenStudy (anonymous):

x is the measure of the smaller angle, thus 17x is the measure of the larger angle. x + 17x = 180 18x = 180 x = 10

OpenStudy (anonymous):

That is, 10 degrees

OpenStudy (anonymous):

x+y=180 x=17y 17y+y=180 18y=180 y=10 x=170